Claudia has received 19 more Christmas cards than Larry.
Larry has received 17 less than Vicki.
Jorge has eight more cards than Lori.
Vicki has received 5 more cards than Jorge.
Jorge and Larry have 44 cards between them.
How many have the 5 received between them?
Thanks for the help I appreciate it. I can't figure this out for the life of me.
Mandie

Answer by stanbon(75887)

Claudia has received 19 more Christmas cards than Larry.
Larry has received 17 less than Vicki.
Jorge has eight more cards than Lori.
Vicki has received 5 more cards than Jorge.
Jorge and Larry have 44 cards between them.
How many have the 5 received between them?
---------------------
C = L + 19
L = V - 17
J = L + 8
V = J + 5
J + L = 44
-------------
Solving:
J + L = 44
(L+8)+L = 44
2L = 36
L = 18
------------------
Then :
C = 18+19 = 37
---------------------
18 = V - 17
V = 35
-----------------
35 = J + 5
J = 30
-----------------
Cheers,
Stan H.